Gaurav is 3 years older than the twins, and five years younger than his sister. the sum of all their ages is 39. how old is each person?
Answer by TimothyLamb(4379)

x = gustav age
y = twins age
z = sister age
---
x = y + 3
x = z - 5
x + 2y + z = 39
---
put the system of linear equations into standard form
---
x - y = 3
x - z = -5
x + 2y + z = 39
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 10
y = 7
z = 15
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---